express the base 16 number 2AD in base 10
Answer by Edwin McCravy(20060)

Hexadecimal is base sixteen.

Numbers spelled out in words are considered the same in any two bases.
The digits are the same in two bases only up through the highest 
digit in the smaller base. 

0 in base ten = 0 in base sixteen
1 in base ten = 1 in base sixteen
2 in base ten = 2 in base sixteen
3 in base ten = 3 in base sixteen
4 in base ten = 4 in base sixteen
5 in base ten = 5 in base sixteen
6 in base ten = 6 in base sixteen
7 in base ten = 7 in base sixteen
8 in base ten = 8 in base sixteen
9 in base ten = 9 in base sixteen
10 in base ten = A in base sixteen
11 in base ten = B in base sixteen
12 in base ten = C in base sixteen
13 in base ten = D in base sixteen
14 in base ten = E in base sixteen
15 in base ten = F in base sixteen

Just as 

276 in base ten means 2×ten² + 7×ten + 6,

2AD in base sixteen means 2×sixteen² + A×sixteen + D

In base ten that's 2×16² + 10×16 + 13 = 2×256+160+13 = 512+160+13 = 685.

Edwin